Hoodoo Voodoo Co. has total assets of $66,750, net working capital of $20,575, owners' equity of $32,435, and long-term debt of $24,265. What is the company's current assets?

Multiple Choice

  • $46,175

  • $32,435

  • $30,625

  • $36,125

  • $56,700
Answer #1

total assets=owners' equity+long-term debt+current liabilities

current liabilities=(66,750-32,435-24,265)=$10050

Net working capital= current assets-current liabilities

current assets=10050+20,575

=$30625